# A Self-Explainable Heterogeneous GNN for Relational Deep Learning

### v1

@@ 4.2 Relation Scoring
Figure 2: Top: First iteration scoring function. For simplicity, we assume the features of the patient nodes to be uninformative, represented by a constant ’$T$’(rue) vector. Top Left: By simultaneous optimization of $w_{2},\ldots,w_{6}$ and $\Theta$ parameters, the loss for relation ’a’ can be brought arbitrarily close to 0 by choosing $\Theta$ so that $Z:=\Theta\cdot T\gg 0$. Top Right:The loss for relation "b" is constant $1/2$ for any choice of the parameters $w_{7},\Theta$. Bottom: Second iteration of the scoring function. With $Z$ as in first iteration we have $\alpha(2,\mathcal{B}_{2})=\alpha(3,\mathcal{B}_{2})=\alpha(5,\mathcal{B}_{3})=\alpha(6,\mathcal{B}_{3})=Z$ and $\alpha(4,\mathcal{B}_{2})=\alpha(4,\mathcal{B}_{3})=2Z$ (details in App. A.2). Bottom Left: With relation "c" optimal parameters give 0 loss. Bottom Right: Relation "d" obtains non-optimal loss (1/2).
@@ 4.3 MPS-GNN
where $\mathcal{N}_{v}^{r_{L-l}}$ are neighbours of node $v$ under relation $r_{L-l}$, $h^{(l)}_{v}$ is the embedding of node $v$ in layer $l$, $h^{(0)}_{v}=x_{v}$ is the feature vector of node $v$, while $W_{0}^{(l)}$, $W_{neigh}^{(l)}$ and $W_{1}^{(l)}$ are learnable parameter vectors. Note that the latter term $W_{1}^{(l)}h^{(0)}_{v}$, which is missing the original MP-GNN [29], represents a skip connection between the input and the $l+1$ layer. This allows the network to access the node attributes at each layer, which is essential for enabling the MPS-GNN to capture node features corresponding to the $\Theta\cdot x_{v}$ terms in the meta-path construction as shown in the ablation study in A.3). As in [29], the definition can be generalized to multiple meta-paths by concatenating the embeddings obtained from each of them.
@@ 4.4 The overall algorithm
Algorithm 1 outlines the whole MPS-GNN procedure for the single meta-path case (in practice, a $K$ beam size is used and multiple meta-paths are learned). The algorithm takes as input a graph ${\cal G}$, the set of available relations ${\cal R}$, an initial set of binary node labels ${\cal Y}$ and a maximal meta-path length $L_{MAX}$. It initializes the targets ${\cal S}$ with a set of singletons (one per labelled node) and their alpha values (collectively indicated by $\mathcal{A}$) to 1. At each iteration, the scoring function identifies the relation minimizing Eq. 3. The algorithm then constructs an MPS-GNN with the current (partial) meta-path, trains it on node labels, and evaluates it using the $F_{1}$ score (computed on a validation set, omitted for brevity), which reflects the meta-path’s performance when embedded in an MPS-GNN, as opposed to its potential informativeness measured by the scoring function.
@@
         $r^{*}\leftarrow\textsc{score}({\cal G},{\cal S},\mathcal{R},\mathcal{A})$
@@
[…]
@@
         $\mathcal{A},{\cal S}\leftarrow\textsc{new-targets}({\cal S},r^{*})$
@@
The algorithm keeps track of the best meta-path prefix together to its $F_{1}$ score, and updates the bags and $\alpha$ values as specified in Equations 5 and 6 for the next relation scoring round. The algorithm ends when the maximum meta-path length $L_{MAX}$ is reached. MPS-GNN scales linearly in the number of relations and nodes, thanks to its incremental construction of meta-paths.
@@ 4.5 MPS-GNN is a self-explainable model
By relying on meta-paths for its predictions, MPS-GNN is a self-explainable model. The scoring function serves as the detector, identifying relevant meta-paths, while the network built using them acts as the classifier. By construction, the network can only access the meta-path induced subgraph, making it strictly sufficient by construction (no changes outside the meta-path induced graph affect the prediction). An analysis of the faithfulness of MPS-GNN’s explanations is provided in Section 5.3.
@@ 5 Experiments
We compared MPS-GNN with approaches that don’t require predefined meta-paths, handle numerous relations, and incorporate node features in learning. The identified competitors include: MLP, to test the sufficiency of target node features alone; GCN [43], a baseline non-relational model; RGCN [44], extending GCN for multi-relational graphs, with distinct parameters for each edge type; HGN [45], a heterogeneous GNN model extending GAT for multiple relations; GTN [46], which transforms input graphs into different meta-path graphs where node representations are learned; Fast-GTN [47], an optimized GTN variant; R-HGNN [48], a relation-aware GNN using cross-relation message passing; and MP-GNN [29], the original meta-path GNN supporting only existentially quantified meta-paths.
We implemented our model using PyTorch Geometric, and used the competitors’ code from their respective papers for comparison. For training MPS-GNN, we used a 70/20/10 split for training, validation, and testing, respectively, and reported the test results for the model selected based on its validation performance. For the sake of comparison with [29], we set the maximum meta-path length to 4 and the beam size to 3. We employed $F_{1}$ as evaluation metric to account for the unbalancing in many of the datasets. The code is freely available at  [^1] https://github.com/francescoferrini/MPS-GNN
5.1 Q1: MPS-GNN consistently identifies the correct meta-path in synthetic scenarios
In order to address the first research question, we designed a sequence of synthetic node classification scenarios where the correct structure to be learnt is known. In each scenario, a node is labelled as positive if it is the starting point of at least $c$ occurrences of a given meta-path of length $l$, and negative otherwise. Crucially, existential quantification of meta-paths (as modelled by MP-GNN [29]) is insufficient here, as nodes which are starting points of less than $c$ meta-path occurrences are labelled as negatives. We designed scenarios of increasing complexity by changing the length of the ground-truth meta-path $l$, the number of occurrences $c$, and the overall number of relations $r$ in the dataset. See Figure 3 for the statistics of the different scenarios (left), and for a prototypical example for $l=2$ and $c=3$ (right).
@@ 5.1 Q1: MPS-GNN consistently identifies the correct meta-path in synthetic scenarios
Table 1 shows the $F_{1}$ score of each model for an increasing complexity of the classification scenario. Results clearly show that this experimental setting is challenging for existing solutions. While the poor performance of MLP, which completely ignores the topological structure, and GCN, which ignores the difference between relations, are expected, solutions specifically conceived for heterogeneous networks also struggle with these datasets. Models like R-HGNN, HGN, GTN, and Fast-GTN, despite accounting for different relations in the graph, are affected by both the imbalance between positive and negative target nodes and the limited number of instances of neighbors of a certain type. RGCN and MP-GNN achieve better performance but are still sub-optimal. The former, like other relational methods, takes into account the diversity of relations in the graph but still uses all of them, thus struggling to single out the relevant portion of the graph. MP-GNN, on the other hand, suffers from its existential quantification assumption, and fails to find the correct meta-path in all scenarios. Conversely, MPS-GNN manages to achieve nearly-optimal performance in all scenarios, substantially outperforming all existing strategies [^2] The residual error for MPS-GNN is due to the fact that despite relying on the correct meta-path, it occasionally leverages spurious instances where the relation sequence is correct but (some of) the node features are not.. These results allow us to answer the first research question in the affirmative.
Table 1: $F_{1}$ scores on the synthetic datasets. Best results are boldfaced. Results are averaged over 5 runs. The full table including standard deviations can be found in Appendix A.4
S1 S2 S3 S4 S5 S6 S7 S8
MLP 0.46 0.44 0.48 0.47 0.44 0.51 0.45 0.47
GCN 0.46 0.46 0.48 0.52 0.44 0.48 0.46 0.48
RGCN 0.78 0.87 0.86 0.81 0.86 0.77 0.91 0.79
R-HGNN 0.50 0.44 0.47 0.47 0.53 0.48 0.46 0.48
HGN 0.45 0.46 0.50 0.46 0.46 0.48 0.45 0.40
GTN 0.46 0.52 0.49 0.48 0.44 0.47 0.49 0.47
Fast-GTN 0.46 0.48 0.51 0.49 0.44 0.47 0.53 0.47
MP-GNN 0.84 0.82 0.85 0.95 0.89 0.79 0.84 0.71
MPS-GNN 0.98 0.98 0.99 0.98 0.99 0.93 0.94 0.95
Our approach is particularly useful for predictive tasks in relational databases with multiple tables, where features for a target entity may involve statistics from related tables. To address the second research question, we thus focused on three relational databases with many tables: EICU, a medical database with 31 tables, where we predict patient stay duration in the eICU; MONDIAL, a geographic database where the task is predicting whether a country’s religion is Christian; and ErgastF1, containing Formula 1 data, where the task is predicting the winner of a race. The databases were transformed into graphs as explained in Section 3; for disconnected components, we enhanced connectivity by clustering rows of auxiliary tables. Additional details for the dataset and the procedure are in the Appendix.
Results
Table 2 presents the $F_{1}$ scores of MPS-GNN and its competitors across three real-world databases, averaged over 5 runs with different seeds. The poor performance of MLP clearly indicates that using target node features only is insufficient for classification. Plain GCN, which treats the graph as homogeneous, performs well only on the EICU dataset, where node degree differences exist between positive and negative nodes. Heterogeneous GNN methods also struggle with these datasets, especially MONDIAL, where most approaches fail to outperform a simple MLP, and only one (HGN) manages to substantially outperform the non-heterogeneous baseline (GCN).
EICU MONDIAL ErgastF1
MLP 0.53 0.52 0.50
GCN 0.89 0.60 0.50
RGCN 0.70 0.53 0.57
R-HGNN 0.61 0.61 0.72
HGN 0.75 0.72 0.70
GTN 0.56 0.38 0.60
Fast-GTN 0.46 0.39 0.60
MP-GNN 0.87 0.36 0.71
MPS-GNN 0.92 0.74 0.83
Table 2: $F_{1}$-score of our method and competitors on the three real-world datasets. Full table with standard deviations can be found in Appendix A.6
MP-GNN does not provide the performance boost that was observed when applied to knowledge graphs [29], confirming our intuition that existential quantification of meta-path is insufficient when dealing with relational databases. On the other hand, MPS-GNN manages to substantially outperform all competitors, thanks to its ability to identify meta-paths that are informative thanks to the statistics that can be computed over their realizations, as shown in the following. It is worth noting that this result is achieved with one/two orders of magnitude fewer parameters than the runner-ups, namely HGN and R-HGNN. See Table 11 in the Appendix for the details. Additionally, Table 12 in the Appendix shows that MPS-GNN has competitive execution times with respect to other heterogeneous GNN approaches, thanks to its ability to focus training on relevant meta-path induced subgraphs.
Identified Meta-Paths
Figure 4: Extracted meta-paths for the three real world datasets.
Finally, in Appendix A.11 we present an experimental evaluation where MPS-GNN is adapted to deal with temporal databases and tasks [31], showing how it outperforms its competitors also in this context. Summing up, these results enable us to confidently answer Q2 in the affirmative.
5.3 Q3: MPS-GNN is a self-explainable method
To address the third research question, we assessed the faithfulness of the extracted meta-paths. The meta-paths identified by the model from the graph are evaluated based on the complementary metrics of sufficiency and necessity. High sufficiency implies that changing the complement to the explanation (leaving the explanation unchanged) should not affect the model’s output. High necessity implies that altering the explanation itself (leaving the complement unchanged) should result in a change in the model’s output. It is easy to show that our approach is inherently sufficient. Indeed, the computational graph of MPS-GNN consists solely of the subgraph containing the occurrences of the identified meta-paths. Necessity, on the other hand, should be experimentally evaluated by altering the computational graph of MPS-GNN.
Table 3: Changes to $F_{1}$ and necessity when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world datasets.
$\mathbf{F_{1}}$ Necessity
Removed (%) 0 25 50 75 0 25 50 75
EICU 0.92 0.82 0.67 0.56 0 0.16 0.18 0.24
MONDIAL 0.74 0.61 0.48 0.43 0 0.13 0.29 0.32
ErgastF1 0.814 0.65 0.61 0.58 0 0.08 0.16 0.26
Table 3 illustrates the effects of removing 25%, 50%, and 75% of the meta-path occurrences in terms of changes in $F_{1}$ and necessity between original and modified graphs [49]. For a visual representation of this table and for the definition of necessity, see Appendix A.7.
In all datasets, there’s a noticeable decline in $F_{1}$ performance and a steep increase in probability difference, suggesting that the learned meta-paths are also necessary. These results clearly indicate the faithfulness of the explanations of our self-explainable method.
6 Conclusion
We introduced a novel approach to identify relevant meta-paths of relations for node classification tasks in heterogeneous graphs with a potentially large number of different relations, notably graphs derived from relational databases. Compared to earlier work, our approach does not require any user supervision, and it can learn meta-paths for predictive features defined by aggregate statistics over meta-path occurrences. Experiments clearly demonstrate advantages over alternative approaches in terms of accuracy, sparsity, and explainability. The approach was presented for binary classification, but it can be straightforwardly extended to multiclass classification using standard one-vs-all strategies. Extending it to regression tasks is an interesting direction for further research.
Appendix A Appendix
A.1 Comparison with MP-GNN
The novelty of our approach compared to MP-GNN lies in our model’s ability to learn meta-paths that are relevant to the target node class, not merely based on their existence but on statistical measures related to their occurrences. As illustrated in the example in Figure 1, MP-GNN would not be able to distinguish between the classes of the two patient nodes, as both have a single occurrence of the correct meta-path. This is due to how the function that iteratively constructs meta-paths in MP-GNN works. Specifically, under the existential quantification assumption, a candidate relation $r$ is informative for the label of a node $v$ f at least one of the neighbors ${\cal N}_{v}^{r}$ of $v$ according to $r$ belongs to the ground-truth meta-path, and $v$ has the right features. To achieve this, MP-GNN selects the optimal relation by identifying the connected node that maximizes information for class label prediction. This represents a significant limitation for MP-GNN, as it cannot differentiate between node classes that are determined by aggregate statistics (e.g. counts) of multiple occurrences of meta-paths.
A.2 Second iteration loss computation
In this Section we show the loss computation and minimization for relation "c" of the example in Figure 2. Note that $Z$ refers to the learned parameters related to the features of the target nodes in the previous iteration as shown in Figure 2 and $\binom{1}{0}$ and $\binom{0}{1}$ represent the features of $prescription$ nodes.
$Z$ $\gg 0$
$F(\mathcal{B}_{2},c)$ $=Z\Theta\binom{0}{1}+2Z\Theta\binom{1}{0}(w_{8}+w_{9})+Z\Theta\binom{1}{0}(w_{10}+w_{11})$
$F(\mathcal{B}_{3},c)$ $=2Z\Theta\binom{1}{0}(w_{8}+w_{9})+Z\Theta\binom{0}{1}(w_{10}+w_{11})+Z\Theta\binom{1}{0}(w_{12})$
$L(c)$ $=\min_{\Theta,w}\sigma\left(F(\mathcal{B}_{3},c)-F(\mathcal{B}_{2},c)\right)$
$=\min_{\Theta,w}\sigma\left(Z\Theta\binom{1}{0}(2w_{8}+2w_{9}+w_{12}-2w_{8}-2w_{9}-w_{10}-w_{11})\right.$
$\quad\left.+Z\Theta\binom{0}{1}(w_{10}+w_{11}-1)\right)$
$=\min_{\Theta,w}\sigma\left(Z\Theta\binom{1}{0}(w_{12}-w_{10}-w_{11})+Z\Theta\binom{0}{1}(w_{10}+w_{11}-1)\right)$
$\Theta\binom{1}{0}$ $\gg 0;\quad\Theta\binom{0}{1}=0;\quad w_{12}=0;\quad w_{10},w_{11}=1$
A.3 Ablation study
In this section, we highlight the importance of the skip connection $W_{1}^{(l)}h_{v}^{(0)}$ considered in equation 7. Table 4 and 5 demonstrate respectively the significance of skip connections in the synthetic and real-world settings. Without considering the initial target node features, the $F_{1}$ score drops significantly, underscoring the critical role these features play.
S1 S2 S3 S4 S5 S6 S7 S8
[…]
EICU MONDIAL ErgastF1
[…]
A.4 Synthetic datasets
In table 6 are presented $F_{1}$-score with standard deviations of MPS-GNN and competitors in the synthetic settings.

### v2

@@ H Refer to caption
4.2.1 Time complexity analysis
The purpose of the scoring function is to iteratively construct the meta-path that best classifies the target nodes, avoiding the need to explore all possible paths. A naive search algorithm that simply tests all possible paths would have a polynomial complexity of $O(|\mathcal{R}|^{L})$, where $|\mathcal{R}|$ is the number of relations in the graph, and $L$ is the maximum length of the meta-path we aim to find. By leveraging the scoring function, the complexity is reduced to linear $O(|\mathcal{R}|\cdot L)$. This improvement is achieved because, at each step, one relationship is added to the meta-path under construction, and the subsequent search builds upon it without reconsidering the other relationships scored in the same iteration.
@@ 4.3 MPS-GNN
[…]
@@ 4.4 The overall algorithm
Algorithm 1 outlines the whole MPS-GNN procedure for the single meta-path case (in practice, a $K$ beam search is used and multiple meta-paths are learned). The algorithm takes as input a graph ${\cal G}$, the set of available relations ${\cal R}$, an initial set of binary node labels ${\cal Y}$ a maximal meta-path length $L_{MAX}$ and a stopping criteria value $\eta$. It initializes the targets ${\cal S}$ with a set of singletons (one per labelled node) and their alpha values (collectively indicated by $\mathcal{A}$) to 1. At each iteration, the scoring function identifies the relation minimizing Eq. 4 and appends it to the meta-path $mp$. If no relation improves the loss by a factor $\eta$, the algorithm stops and returns the meta-path $mp^{*}$ constructed so far. The algorithm then evaluates $mp$ by training MPS-GNN on node labels, and tests it using the $F_{1}$ score (computed on a validation set, omitted for brevity), which reflects the meta-path’s performance when embedded in an MPS-GNN, as opposed to its potential informativeness measured by the scoring function. For training MPS-GNN, node embeddings are updated using 7.
@@
         $r^{*}\leftarrow argmin_{r}L(r)$triangleright Eq. 4
         if $\min_{r\in\mathcal{R}}L(r)\geq\eta L_{\text{init}}(r)$ then
              return $mp^{*}$
         end if
@@
         $gnn\leftarrow\textsc{train}(\textsc{MPS-GNN}(mp),{\cal G},{\cal Y})$
         $F_{1}\leftarrow\textsc{test}(gnn)$
@@
         $\mathcal{A},{\cal S}\leftarrow\textsc{new-targets}({\cal S},r^{*})$triangleright Eq. 5, 6
@@
The algorithm keeps track of the best meta-path prefix together to its $F_{1}$ score, and creates new target bags and $\alpha$ values as specified in Equations 5 and 6 for the next relation scoring round. The algorithm ends when the maximum meta-path length $L_{MAX}$ is reached The algorithm is described for simplicity in the context of a single meta-path and without the additional stopping criteria described in section 4.2. In practice, however, the implementation employs a beam search over the meta-path space, selecting the top $K$ best-scoring meta-paths concatenating their embedding for the final node representation as detailed at the end of section 4.3. MPS-GNN scales linearly in the number of relations and nodes, thanks to its incremental construction of meta-paths. The value chosen for $L_{\text{MAX}}$ in the experiment was 4, while $\eta$ was set to 0.7. The method is however quite robust to variations in $\eta$, as the identified meta-paths remained unchanged for a wide range of values in our experiments.
@@ 4.5 MPS-GNN is a self-explainable model
Self-explainable GNNs (Kakkad et al., 2023; Christiansen et al., 2023) are a class of GNN models that aims to achieve explainability by-design. At a high level, these models can be seen as composed of two modules: a detector that extracts a class-discriminative subgraph, and a classifier that outputs a prediction based on the extracted subgraph. By relying on meta-paths for its predictions, MPS-GNN is a self-explainable GNN model. The scoring function serves as the detector, identifying relevant meta-paths, while the network built using them acts as the classifier. By construction, the network can only access the meta-path induced subgraph, making it strictly sufficient by construction (no changes outside the meta-path induced graph affect the prediction). An analysis of the faithfulness of MPS-GNN’s explanations is provided in Section 5.3. While most approaches for GNN explainability focus on the topological aspect, by identifying (hard or soft) subgraphs as explanations, determining which node features contribute to the prediction is also relevant from an interpretability perspective. While interpretability at the node feature level could be encouraged by introducing sparsifying norms for the learnable parameter vectors in Eq. 7, guaranteeing a fully transparent processing of node features in the layerwise node embedding computation is beyond the scope of current GNN-based architectures.
4.6 Comparison with MP-GNN
The novelty of our approach compared to MP-GNN lies in our model’s ability to learn meta-paths that are relevant to the target node class, not merely based on their existence but on statistical measures related to their occurrences. In practice the key difference between MPS-GNN and MP-GNN lies in the way in which relations are scored. Specifically MP-GNN predicts the class label $y$ of a node $v$, in the first iteration, with $\tilde{y}_{v}^{r}=\Theta^{T}x_{v}\max_{u\in{\cal N}_{v}^{r}}w_{u}$ where the $max$ aggregation of the neighbours is used, indicating that a candidate relation $r$ is informative for the label of a node $v$ if at least one of the neighbors ${\cal N}_{v}^{r}$ of $v$ according to $r$ belongs to the ground-truth meta-path, and $v$ has the right features. To leverage meta-path occurrences, MPS-GNN employs a $sum$ aggregation strategy, as detailed in Eq. 2, enabling the counting of their occurrences.
In subsequent iterations, MP-GNN also employs a bag creation process, where positive bags are formed by including neighbors of positive nodes that are predicted to be positive at least once across multiple prediction procedures. This formulation ensures that the node responsible for the positive label, and thus aligned with the correct meta-path, remains in a positive bag. Unlike this approach, however, we must ensure that all neighbor nodes involved in multiple occurrences of the $correct$ meta-path are included in a positive bag. As detailed in eq. 5, the neighbors of positive nodes are therefore placed in a positive bag without the need of any additional prediction step.
The enhancements introduced in MPS-GNN, compared to MP-GNN, enable it to handle predictions over relational databases where the class label may depend on statistical measures derived from meta-path occurrences.
@@ 5 Experiments
We compared MPS-GNN with approaches that don’t require predefined meta-paths, handle numerous relations, and incorporate node features in learning. The identified competitors include: MLP, to test the sufficiency of target node features alone; GCN (Kipf & Welling, 2016), a baseline non-relational model; RGCN (Schlichtkrull et al., 2017), extending GCN for multi-relational graphs, with distinct parameters for each edge type; HGN (Lv et al., 2021a), a heterogeneous GNN model extending GAT for multiple relations; GTN (Yun et al., 2019a), which transforms input graphs into different meta-path graphs where node representations are learned; Fast-GTN (Yun et al., 2022b), an optimized GTN variant; R-HGNN (Yu et al., 2021), a relation-aware GNN using cross-relation message passing; and MP-GNN (Ferrini et al., 2024), the original meta-path GNN supporting only existentially quantified meta-paths.
We implemented our model using PyTorch Geometric, and used the competitors’ code from their respective papers for comparison. For training MPS-GNN, we used a 70/20/10 split for training, validation, and testing, respectively, and reported the test results for the model selected based on its validation performance. For the sake of comparison with Ferrini et al. (2024), we set the maximum meta-path length to 4 and the beam size to 3. We employed $F_{1}$ as evaluation metric to account for the unbalancing in many of the datasets. The code is freely available at  [^2] https://github.com/francescoferrini/MPS-GNN. Hyperparameters of competitors and MPS-GNN can be found in Table 8 in the Appendix.
5.1 Q1: MPS-GNN consistently identifies the correct meta-path in count based synthetic scenarios
In order to address the first research question, we designed a sequence of synthetic node classification scenarios where the correct structure to be learnt is known. In each scenario, a node is labelled as positive if it is the starting point of at least $c$ occurrences of a given meta-path of length $l$, and negative otherwise. Crucially, existential quantification of meta-paths (as modelled by MP-GNN (Ferrini et al., 2024)) is insufficient here, as nodes which are starting points of less than $c$ meta-path occurrences are labelled as negatives. We designed scenarios of increasing complexity by changing the length of the ground-truth meta-path $l$, the number of occurrences $c$, and the overall number of relations $r$ in the dataset. See Figure 6 for the statistics of the different scenarios (left), and for a prototypical example for $l=2$ and $c=3$ (right).
Correct Pattern
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
S1 S2 S3 S4 S5 S6 S7 S8
$|R|$ 5 5 5 5 10 10 10 10
$c$ 2 2 3 4 2 2 3 4
$l$ 2 3 3 2 2 3 3 2
@@ 5.1 Q1: MPS-GNN consistently identifies the correct meta-path in count based synthetic scenarios
Table 1 shows the $F_{1}$ score of each model for an increasing complexity of the classification scenario. Results clearly show that this experimental setting is challenging for existing solutions. While the poor performance of MLP, which completely ignores the topological structure, and GCN, which ignores the difference between relations, are expected, solutions specifically conceived for heterogeneous networks also struggle with these datasets. Models like R-HGNN, HGN, GTN, and Fast-GTN, despite accounting for different relations in the graph, are affected by both the imbalance between positive and negative target nodes and the limited number of instances of neighbors of a certain type. RGCN and MP-GNN achieve better performance but are still sub-optimal. The former, like other relational methods, takes into account the diversity of relations in the graph but still uses all of them, thus struggling to single out the relevant portion of the graph. MP-GNN, on the other hand, suffers from its existential quantification assumption, and fails to find the correct meta-path in all scenarios. Conversely, MPS-GNN manages to achieve nearly-optimal performance in all scenarios, substantially outperforming all existing strategies [^3] The residual error for MPS-GNN is due to the fact that despite relying on the correct meta-path, it occasionally leverages spurious instances where the relation sequence is correct but (some of) the node features are not.. Note that the lookahead capabilities of the scoring function are crucial to the effectiveness of MPS-GNN. Appendix A.7 shows how replacing the scoring function with a simple greedy approach leads to failure in learning the correct meta-path. These results allow us to answer the first research question in the affirmative.
Table 1: $F_{1}$ metric with standard deviations for synthetic datasets